OpenStudy (anonymous):

Determine the solutions to the quadratic equation -20x2 - 20x + 40 = 0.

OpenStudy (anonymous):

-20(x^2+x-2)=0 x^2+x-2=0 x^2+2x-x-2=0 x(x+2)-(x+2)=0 (x-1)(x+2)=0 x=1,x=-2